OpenStudy (anonymous):

A store received $823 from the sale of 5 tape recorders and 7 radios. If the receipts from the tape recorders exceeded the receipts from the radios by $137, what was the cost of a tape recorder? $49 $68 $84 $96

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

So: x - price of all radios x+137 - price of all tape recorders than: x+137+x=823 2x=686 x=343 => price of 7 radios = 343 => price of 5 tape recorders = 823-343 = 480 than 1 tape recorder's price is 480/5 = 96. I think this way:)

OpenStudy (anonymous):

TAPE RECORDER=TP RADIO=R 5TP+7R=823 5TP-7R=137 10TP= 960 TP=960/10=96